OpenStudy (anonymous):

you want to evaluate the integral xsin(5x^2)dx using substituiton. which of the following is the best choice for the substitution variable u? A. u=3x-x^3 B. u=4x^2 C.(3x-x^3)^5 D. x^2(3x-x^3)^5

OpenStudy (anonymous):

Do you have any questions on the method of substitution?

OpenStudy (anonymous):

u= xsin(5x^2) u= sin(5x^2) u= x u= 5x^2 i got u= sin(5x^2).....

OpenStudy (anonymous):

The derivative of 5x^2 is 10x, so try sin(u) du.

OpenStudy (anonymous):

Let u=5x^2 then see what you need to do to get from dx to du.